The initial kinetic energy imparted to a 0.022 bullet is 1300 J. Find the range of this projectile when it is fired at an angle such that the range is equals the maximum height attained.


1
Expert's answer
2021-05-30T13:30:32-0400

"l_{max}=v^2\sin(2\alpha)/g"


"h_{max}=v^2\sin^2\alpha/(2g)"


"v^2\sin^2\alpha/(2g)=v^2\sin(2\alpha)/g\to\alpha=75.96°"


"mv^2/2=1300\to v=\sqrt{2600/0.022}=344\ (m/s)". So,


"l_{max}=v^2\sin(2\alpha)/g=344^2\cdot\sin(2\cdot75.96)/9.8=5684\ (m)" . Answer